# Visualized Text-to-Image Retrieval

### v1

@@ Appendix C Implementation Details
To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in it. The prompts for three benchmarking datasets are shown in Figure 3, Figure 4, and Figure 5, respectively. Then, we wrap the rephrased query with a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation. We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high". For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a signle instruction already results in images generated with a high level of diversity. Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
@@
We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (Wei et al., 2022). Concretely, the model is asked to always extract visual information, perform reasoning, and conclude its reasoning with self-verification. We show the detailed prompt in Figure 6 and Figure 7.
@@
For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important feature that the query is seeking. At the early stage of the prompt, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction. Therefore, we report the results with the same prompt for the final version for the paper.
VQA Evaluation
For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025), as shown in Figure 8. Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering. We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
Dataset Question Ground Truth Baseline Generated Image VisRet
@@
Mexican grass-carrying wasp visiting a purple flower [Uncaptioned image] Rank:45, NDCG@10: 0.00 [Uncaptioned image] Rank:4, NDCG@10: 0.39
@@
Figure 5: Prompt for instructing an LLM to generate the T2I generation instruction for INQUIRE-Rerank questions.

### v2

@@ Appendix B Implementation Details
To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in it. The prompts for three benchmarking datasets are shown in Figure˜3, Figure˜4, and Figure˜5, respectively. Then, we wrap the rephrased query in a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation. We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high". For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a single instruction already results in images generated with a high level of diversity. Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
@@
We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (Wei et al., 2022). Concretely, the model is asked to always extract visual information, perform reasoning, and conclude its reasoning with self-verification. We show the detailed prompt in Figure˜6 and Figure˜7.
VQA Evaluation
For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025b), as shown in Figure˜8. Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering. We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
@@
For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important feature that the query is seeking. At the early stage of the project, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction. Therefore, we report the results with the same prompt for the final version for the paper. For the Corpus Captioning baseline, we use the Salesforce/blip-image-captioning-large model accessed at huggingface. We follow the recommended parameters to run inference on local GPUs, except setting the minimum output length to 10 tokens to encourage a more detailed caption. For the VISA Reranking baseline, we followed the original paper’s public implementation [^6] https://github.com/XLearning-SCU/2025-ICML-VISA, using the sample prompt except changing the model to Qwen2.5-VL-72B-Instruct which we empirically find better performance. For the Captioning Reranking baseline, we simplify the VISA pipeline by directly prompting for a caption for each of the top retrieved images using the following prompt:
"Write a caption that describes the appearance and posture of the main subject. The caption should be concise but informative so that it can be used to retrieve the images from images of the same subject in different postures, states, or environments."
Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
@@
INQUIRE Mexican grass-carrying wasp visiting a purple flower [Uncaptioned image] Rank:45, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.39
@@
COCO A den with a large coffee table, couch and a television. [Uncaptioned image] Rank:61, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.3937
A computer workstation with a laptop and a desktop computer. [Uncaptioned image] Rank:37, nDCG@10: 0.00 [Uncaptioned image] Rank:1, nDCG@10: 1.0
A residential bathroom with sink, tub, and toilet setting in it. [Uncaptioned image] Rank:54, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.39
@@
Figure 5: Prompt for instructing an LLM to generate the T2I generation instruction for INQUIRE-Rerank-Hard and COCO-Hard questions.
@@
Appendix C Further Qualitative Examples
In Table˜8, we present additional qualitative examples on the four evaluation datasets. Consistent with quantitative findings, VisRet reduces the embedding’s workload by expressing the important details in a visualization close to the retrieval target, resulting in much better performance in terms of ground truth rank and nDCG@10.